OpenStudy (anonymous):

Solve for x: 2(4x + 12) + 4x = -12 A. -3 B. -2 C. -1 D. 1

OpenStudy (anonymous):

2(4x+12)+4x=-12 8x+24+4x=-12 12x+24=-12 12x=-36 x=-3

OpenStudy (anonymous):

It was A